# JWST spectroscopy of a blue binary cold classical Kuiper belt object

### v1

@@ Abstract
We present observations of two binary systems within the cold classical region of the Kuiper belt—2001 XR254 and 2016 BP81—obtained with the JWST Near-Infrared Spectrograph. The measured reflectance spectrum of 2001 XR254 is characteristic of the red cold classicals, with strong features due to carbon dioxide, carbon monoxide, and methanol ices. In contrast, 2016 BP81 is a blue binary, with a water-ice-rich surface composition. The two components of the 2016 BP81 binary display identical spectral profiles, consistent with coeval formation from gravitational collapse. Through qualitative and quantitative comparisons of water-ice-rich Kuiper belt objects observed with JWST, we identify a small subclass, including 2016 BP81, that appears to differ in systematic ways from the rest of the population. The relatively deep carbon dioxide ice absorption bands and enhanced signatures of aliphatic organics suggest that objects within this subclass may have originated in a distinct formation environment from the other water-ice-rich Kuiper belt objects. The implications of our findings are discussed within the context of recent models of Kuiper belt formation and evolution.
@@ 1 Introduction
The vast population of icy bodies in the Kuiper belt traces the initial compositional landscape and subsequent dynamical processes in the outer protoplanetary disk and holds the key to understanding many fundamental aspects of early solar system history. As the detailed orbital architecture of the Kuiper belt has come into focus in the past two decades, a surprisingly complex dynamical picture has emerged (e.g., B. Gladman et al., 2008; J. Kavelaars et al., 2008; M. E. Brown, 2012; J. J. Kavelaars et al., 2020), which has in turn motivated new theories of solar system evolution. Recent models attribute the broad eccentricity and inclination distribution of Kuiper belt objects (KBOs) to outward migration of Neptune that occurred shortly after the end of planet formation (A. Morbidelli & D. Nesvorný, 2020; B. Gladman & K. Volk, 2021). Numerical simulations have successfully reproduced most dynamical features of the present-day Kuiper belt by invoking a gradual expansion of Neptune’s orbit, which led to the disruption of the primordial disk of icy planetesimals that was situated beyond 30 au (e.g., H. F. Levison et al., 2008; D. Nesvorný & A. Morbidelli, 2012; D. Nesvorný & D. Vokrouhlický, 2016).
@@
Contemporaneously, a concerted effort has been levied toward characterizing the surface properties of KBOs. Ground- and space-based multiband photometry and low-resolution spectroscopy uncovered a stark color bimodality at both visible and near-infrared wavelengths among objects fainter than an absolute magnitude of mag, i.e., smaller than 300–400 km (e.g., W. C. Fraser & M. E. Brown, 2012; W. C. Fraser et al., 2023; N. Peixinho et al., 2012, 2015; S. C. Tegler et al., 2016; I. Wong & M. E. Brown, 2017; M. E. Schwamb et al., 2019). This bimodality separates KBOs into two subpopulations that we refer to here as blue and red.111Other works in the literature apply different binary color designations, e.g., gray/red, red/very-red. The color bimodality suggests that a sharp gradient in surface composition must have existed somewhere within the primordial transneptunian disk, possibly caused by differential sublimation and irradiation chemistry that occurred on two sides of a volatile ice retention line (E. L. Schaller & M. E. Brown, 2007; M. E. Brown et al., 2011; I. Wong & M. E. Brown, 2016).
@@
The advent of JWST has provided the means to place the attested diversity of KBO colors within a more concrete compositional framework. Near-infrared observations of dozens of KBOs in the range mag have revealed three distinct compositional classes (M. N. De Prá et al., 2025; J. Licandro et al., 2025; N. Pinilla-Alonso et al., 2025): (1) objects with deep water-ice (H2O) bands and blue visible spectral slopes that also show minor absorptions due to carbon dioxide (CO2) ice; (2) CO2-rich objects, with colors consistent with the red visible color subpopulation, that display an exceptionally deep 4.27-µm CO2 primary band, evidence of carbon monoxide (CO) ice, and strong organic features produced from the irradiation of aliphatic hydrocarbons (e.g., methane, ethane); (3) a separate subclass of red-colored KBOs with a distinct near-infrared continuum shape that additionally show characteristic features of methanol (CH3OH).222These three compositional classes are designated by the respective monikers ‘bowl’, ‘double-dip’, and ‘cliff’ in N. Pinilla-Alonso et al. (2025). Extending the aforementioned volatile ice retention hypothesis to the three compositional classes implies the existence of two compositional gradients within the primordial Kuiper belt region. Based on the observed compositional inventories, the volatile-depleted H2O-rich KBOs are expected to have formed in the innermost region of the disk, with the CO2-rich and CH3OH-rich classes accreting at progressively larger heliocentric distances (N. Pinilla-Alonso et al., 2025).
@@
The so-called cold classicals—KBOs on relatively low-inclination, low-eccentricity orbits that lie between the 3:2 and 2:1 mean motion resonances with Neptune at 42–47 au—challenge our current understanding of Kuiper belt formation and evolution. This population is notable for containing a high fraction (30%) of near-equal-mass binaries, including many with wide separations (K. Noll et al., 2020). Dynamical studies have demonstrated that a majority of these binary systems would have been lost had they been scattered into their present-day orbits during the outward migration of Neptune (A. H. Parker & J. J. Kavelaars, 2010; D. Nesvorný & D. Vokrouhlický, 2019). Instead, cold classicals may have formed in situ and therefore comprise a unique representative collection of planetesimals from the outermost reaches of the protoplanetary disk (D. Nesvorný et al., 2022).
@@
Within the aforementioned compositional gradient model, one would expect in situ formation of cold classicals to exclusively produce red objects. While red objects indeed dominate the cold classical population, observations within the last decade have revealed a significant admixture of bluer objects in this region; moreover, these bluer objects are predominantly binaries (W. C. Fraser et al., 2017, 2021). These blue binaries are unlikely to have formed within the inner portion of the primordial transneptunian planetesimal disk, given the fragility of widely-separated binaries to significant dynamical excitation. Instead, it has been proposed that all blue binaries formed as binary pairs immediately inward of the present-day cold classical region, between 38 and 42 au; subsequent slow migration of Neptune’s 2:1 mean motion resonance could have gently pushed these objects out into their current orbits, consistent with the somewhat inflated inclination distribution of blue binaries when compared to the redder cold classicals (W. C. Fraser et al., 2017, 2021).
@@
However, recent attempts to simulate this evolutionary pathway have reached an impasse (e.g., D. Nesvorný et al., 2020, 2022). While a color transition immediately inward of the current cold classical region readily produces the blue binaries, it also predicts a significantly larger proportion of bluer objects throughout the inner transneptunian planetesimal disk, particularly below 30 au, where the vast majority of the primordial mass is expected to be concentrated. As a result of Neptune’s migration, a large number of blue singletons would have been scattered outward and implanted into the cold classical population, inconsistent with the observed preponderance of binaries. A drastic reformulation of the initial conditions in the transneptunian region and the specific trajectory of Neptune’s migration would be needed to resolve this quandary. An alternative explanation posits that all cold classicals (including the blue binaries) formed in situ, with the color distinction being a consequence of the timing of formation relative to the dispersion of the gas disk (D. Nesvorný et al., 2022).
@@
2001 XR254 | 42.9 | 0.035 | 1.2 | 172, 140 | 9310 | 2024 Nov 21 05:13:26 | 43.79 | 43.45 | 1.23 | 22.3 | 2976 | ||
2016 BP81 | 43.7 | 0.074 | 4.2 | 188, 170 | 11300 | 2025 Jan 18 15:52:09 | 43.29 | 43.27 | 1.31 | 22.6 | 3793 | pointing offset | |
2003 HG57 | 43.8 | 0.038 | 2.1 | 156, 156 | 13200 | 2025 Feb 15 15:23:07 | 42.45 | 42.31 | 1.33 | 22.8 | 5019 | target missed |
Note. : orbital semimajor axis; : orbital eccentricity; : orbital inclinationl , : diameters of the primary and secondary components; : mutual separation distance; : visible spectral slope; : observation start time; : heliocentric distance at time of observation; : distance from JWST at time of observation; : phase angle of observation; : visible apparent magnitude at time of observation; : total exposure time across all four dithers. The orbital parameters and viewing geometry are computed by JPL Horizons. The visible spectral slope measurements are tabulated in W. C. Fraser et al. (2021). The sizes and mutual separation of the binary components are taken from D. Nesvorný et al. (2022).
@@ 2.1 Observation Design
The observations were carried out using the integral field unit (IFU) on the Near-Infrared Spectrograph (NIRSpec; P. Jakobsen et al. 2022; T. Böker et al. 2023). A set of four dithered exposures was collected for each target using the PRISM grating, which covers the 0.6–5.3 µmwavelength range at an average spectral resolution of . To reduce detector readnoise, all exposures were read out with the NRSIRS2RAPID method (S. H. Moseley et al., 2010; B. J. Rauscher et al., 2012). Due to the large uncertainties in the binary orbit solutions for our targets, no timing constraints were stipulated to schedule the observations to coincide with maximum component separation.
@@ 2.2 Target Selection
When selecting targets for observation as part of our Cycle 3 JWST program, we consulted the list of cold classicals with visible spectral slope measurements published in W. C. Fraser et al. (2021). Restricting our search to confirmed binaries and adopting the spectral slope value of nm as the threshold dividing the blue and red color subpopulations, we chose the three brightest blue binaries according to their minimum apparent magnitude during the Cycle 3 observing windows. The selected targets were 2001 XR254, 2003 HG57, and 2016 BP81, which have measured spectral slopes of , , and , respectively, in units of %/100 nm. Table 1 provides additional information about the targets, including component diameters, mutual separations, and the details of our observations.
@@ 2.3 Data Reduction
Data processing and spectral extraction were handled by the jwstspec pipeline333https://github.com/ianyuwong/jwstspec—a dedicated tool for customized end-to-end reduction of JWST spectroscopic datasets that was developed primarily for observations of small bodies in the Solar System. This pipeline has been employed in numerous previous analyses of NIRSpec observations (e.g., J. P. Emery et al., 2024; W. M. Grundy et al., 2024; I. Wong et al., 2024), and a full description of the overall pipeline workflow can be found in those publications. Beginning with the uncalibrated data, jwstspec passed through the first two stages of Version 1.17.1 of the official JWST pipeline (H. Bushouse et al., 2025), with necessary reference files drawn from context jwst_1322.pmap of the JWST Calibration Reference Data System (CRDS). All pipeline parameters were left at their default values, except for the nsclean step in the second stage, which was manually activated to correct for residual readnoise on the detector. The result of the pipeline processing was bias-corrected, flat-fielded, spatially-rectified, wavelength-calibrated, and flux-calibrated IFU data cubes for each dithered exposure.
@@
The jwstspec pipeline offers several different spectral extraction techniques. For the results presented in this paper, we utilized the empirical PSF fitting method, which is optimized for faint point-source targets and yields substantial improvements in precision over standard circular aperture extraction (e.g., I. Wong et al., 2024; N. Pinilla-Alonso et al., 2025). In short, a moving median window of a specified width is applied along the wavelength axis to generate background-subtracted template PSFs, which are then normalized and fit to the central wavelength slices within the windows to obtain the best-fit multiplicative scaling factors for the templates. The fluxes are then measured from the scaled template PSFs within a user-defined circular aperture. See the referenced papers for full details on this method. For our observations, we set the background region as all pixels outside of a box centered on the measured centroid. The moving median window width was set to 51 (i.e., from to along the wavelength axis), though adjusting this value by up to did not substantively change the resultant spectra. Uncertainties on the fluxes were estimated by weighting the error arrays with the peak-normalized PSF template model and adding in quadrature the uncertainties of all pixels within the extraction aperture.
@@
The spectra extracted from the set of four dithered exposures were passed through a 21-pixel-wide moving median filter to remove outliers. The combined irradiance spectrum of each target was obtained by a simple mean of the four individual dithered spectra. Due to low instrument throughput and poor data quality at the extreme ends of the wavelength range, we trimmed the spectra to 0.7–5.1 um.
@@ 2.4 Accounting for Sensitivity Variations
We used NIRSpec observations of the G-type solar analog star SNAP-2 when deriving the reflectance spectra of our targets. By dividing the irradiance spectrum of an object by that of the star, processed in an analogous manner and extracted from an aperture with the same equivalent encircled flux, we self-consistently corrected for the wavelength-dependent flux losses outside of the extraction apertures, while also removing any common-mode instrumental systematics that are present in the data. In the case of the blended binary 2001 XR254, our choice of the 4-pixel-radius aperture is roughly equivalent to two 3-pixel-radius apertures centered on the two components in terms of encircled flux, and we therefore used a 3-pixel-radius aperture when extracting the stellar spectrum. For 2016 BP81, the stellar spectrum was extracted with the same 2-pixel-radius aperture as the target spectrum.
The offset of 2016 BP81 relative to the nominal pointing region in the center of the FOV is almost , which is larger than the range subtended by the 4-point dither pattern. This raises the possibility of systematic sensitivity variations across the FOV that could bias the shape of the resultant reflectance spectrum when incorporating stellar observations in which the source was centered in a different region of the FOV. To investigate this issue, we considered two observations of SNAP-2, one that employed a 20-point dither pattern (Program #1128, Observation 2) and another that used the 4-point nod pattern (Program #4498, Observation 19).444For more on NIRSpec IFU dither and nod patterns, see https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-operations/nirspec-dithers-and-nods/nirspec-ifu-dither-and-nod-patterns. In the former case, the spread in pointing positions is similar to the 4-point dither pattern that was selected for our KBO targets, while in the latter, the inter-nod separations are much larger, with the star placed at the vertices of a box centered on the middle of the FOV.
Figure 2 shows the irradiance spectra of SNAP-2 from the two observations, both extracted using a 3-pixel-radius aperture. Also plotted is the CALSPEC model spectrum of SNAP-2 (R. C. Bohlin et al., 2014). The two NIRSpec spectra have largely equivalent shapes, except at wavelengths shorter than about 1.4 µm: here, the spectrum measured from the 4-point nod observation drops off significantly, indicating a relative decrease in detector sensitivity toward the edges of the FOV. Meanwhile, the spectrum derived from the 20-point dither pattern is in good agreement with the CALSPEC model. It follows that the choice of one or the other SNAP-2 spectrum affects the continuum shape of the resultant reflectance spectrum, specifically the short-wavelength spectral slope.

### v2

@@ Abstract
We present observations of two binary systems within the cold classical region of the Kuiper belt—2001 XR254 and 2016 BP81—obtained with the JWST Near-Infrared Spectrograph. The measured reflectance spectrum of 2001 XR254 is characteristic of the red cold classicals, with strong features due to carbon dioxide, carbon monoxide, and methanol ices. In contrast, 2016 BP81 is a blue binary, with a spectrum that shows prominent water-ice absorption bands. The two components of the 2016 BP81 binary display identical spectral profiles, consistent with coeval formation from gravitational collapse. Through qualitative and quantitative comparisons of Kuiper belt objects (KBOs) with similar spectral profiles dominated by water-ice features, we identify a small cluster, including 2016 BP81, that differs in systematic ways from the rest of the population. The relatively deep carbon dioxide ice absorption bands, different water-ice band-depth ratios, and possibly enhanced signatures of aliphatic organics suggest that objects in this cluster may have originated in a distinct formation environment from the other water-type KBOs. The implications of our findings are discussed within the context of recent models of Kuiper belt formation and evolution.
@@ 1 Introduction
The vast population of icy bodies in the Kuiper belt traces the initial compositional landscape and subsequent dynamical processes in the outer protoplanetary disk and holds the key to understanding many fundamental aspects of early solar system history. As the detailed orbital architecture of the Kuiper belt has come into focus in the past two decades, a surprisingly complex dynamical picture has emerged (e.g., gladman2008; kavelaars2008; brown2012; kavelaars2020), which has in turn motivated new theories of solar system evolution. Recent models attribute the broad eccentricity and inclination distribution of Kuiper belt objects (KBOs) to outward migration of Neptune that occurred shortly after the end of planet formation (morbidelli2020; gladman2021). Numerical simulations have successfully reproduced most dynamical features of the present-day Kuiper belt by invoking a gradual expansion of Neptune’s orbit, which led to the disruption of the primordial disk of icy planetesimals that was situated beyond 30 au (e.g., levison2008; nesvorny2012; nesvorny2016).
@@
Contemporaneously, a concerted effort has been levied toward characterizing the surface properties of KBOs. Ground- and space-based multiband photometry and low-resolution spectroscopy uncovered a stark color bimodality at both visible and near-infrared wavelengths among objects fainter than an absolute magnitude of mag, i.e., smaller than 300–400 km (e.g., fraser2012; peixinho2012; peixinho2015; tegler2016; wong2017; schwamb2019; fraser2023). This bimodality separates KBOs into two subpopulations that we refer to here as less red and red.111Other works in the literature apply different binary color designations, e.g., gray/red and red/very red. The color bimodality suggests that a sharp gradient in surface composition must have existed somewhere within the primordial trans-Neptunian disk, possibly caused by differential sublimation and irradiation chemistry that occurred on two sides of a volatile ice retention line (schaller2007; brown2011; wong2016).
@@
The advent of JWST has provided the means to place the attested diversity of KBO colors within a more concrete compositional framework. Near-infrared observations of dozens of KBOs in the range mag have revealed three distinct compositional classes (depra2025; licandro2025; pinillaalonso2025): (1) objects with deep water-ice (H2O) bands and less-red visible spectral slopes that also show minor absorptions due to carbon dioxide (CO2) ice; (2) objects with colors consistent with the red visible color subpopulation that display an exceptionally deep 4.27 m CO2 primary band, evidence of carbon monoxide (CO) ice, and strong organic features produced from the irradiation of aliphatic hydrocarbons (e.g., methane, ethane); (3) a separate class of red-colored KBOs with a distinct near-infrared continuum shape that additionally show characteristic features of methanol (CH3OH). In the following, we adopt the recent taxonomic nomenclature proposed in holler2025rnaas and refer to these three compositional classes as H2O-type, CO2-type, and organics-type, respectively.222These three compositional classes are designated by the respective monikers “bowl,” ‘double-dip,” and “cliff” in pinillaalonso2025. Extending the aforementioned volatile ice retention hypothesis to the three types of KBOs implies the existence of two compositional gradients in the surface chemistry of planetesimals within the primordial Kuiper belt region. Based on the observed compositional inventories, the H2O-type KBOs, which show the fewest volatile ice signatures on their surfaces, are expected to have formed in the innermost region of the disk, with the CO2- and organics-type objects accreting at progressively larger heliocentric distances (pinillaalonso2025).
@@
The so-called cold classicals—KBOs on relatively low-inclination, low-eccentricity orbits that lie between the 3:2 and 2:1 mean motion resonances with Neptune at 42–47 au—challenge our current understanding of Kuiper belt formation and evolution. This population is notable for containing a high fraction (30%) of near-equal-mass binaries, including many with wide separations (noll2020). Dynamical studies have demonstrated that a majority of these binary systems would have been lost had they been scattered into their present-day orbits during the outward migration of Neptune (parker2010; nesvorny2019a). Instead, cold classicals may have formed in situ and therefore comprise a unique representative collection of planetesimals from the outermost reaches of the protoplanetary disk (nesvorny2022).
@@
Within the aforementioned compositional gradient model, one would expect in situ formation of cold classicals to exclusively produce red objects. While red objects indeed dominate the cold classical population, observations within the last decade have revealed a significant admixture of objects in this region with bluer visible spectral slopes; moreover, these bluer objects are predominantly binaries (fraser2017; fraser2021). These objects are commonly referred to as blue binaries in the literature, though we note that their colors are broadly consistent with the less-red KBO subpopulation. The blue binaries are unlikely to have formed within the inner portion of the primordial trans-Neptunian planetesimal disk, given the fragility of widely separated binaries to significant dynamical excitation. Instead, it has been proposed that all blue binaries formed as binary pairs immediately inward of the present-day cold classical region, between 38 and 42 au; subsequent slow migration of Neptune’s 2:1 mean motion resonance could have gently pushed these objects out into their current orbits, consistent with the somewhat inflated inclination distribution of blue binaries when compared to the redder cold classicals (fraser2017; fraser2021).
@@
However, recent attempts to simulate this evolutionary pathway have reached an impasse (e.g., nesvorny2020; nesvorny2022). While a color transition immediately inward of the current cold classical region readily produces the blue binaries, it also predicts a significantly larger proportion of less-red objects throughout the inner trans-Neptunian planetesimal disk, particularly below 30 au, where the vast majority of the primordial mass is expected to be concentrated. As a result of Neptune’s migration, a large number of less-red singletons would have been scattered outward and implanted into the cold classical population, inconsistent with the observed preponderance of binaries. A drastic reformulation of the initial conditions in the trans-Neptunian region and the specific trajectory of Neptune’s migration would be needed to resolve this quandary. An alternative explanation posits that all cold classicals (including the blue binaries) formed in situ, with the color distinction being a consequence of the timing of formation relative to the dispersion of the gas disk (nesvorny2022).
@@
| 2001 XR254 | 42.9 | 0.035 | 1.2 | 172, 140 | 9310 | 2024 Nov 21 05:13:26 | 43.79 | 43.45 | 1.23 | 22.3 | 2976 | ||
| 2016 BP81 | 43.7 | 0.074 | 4.2 | 188, 170 | 11300 | 2025 Jan 18 15:52:09 | 43.29 | 43.27 | 1.31 | 22.6 | 3793 | pointing offset | |
| 2003 HG57 | 43.8 | 0.038 | 2.1 | 156, 156 | 13200 | 2025 Feb 15 15:23:07 | 42.45 | 42.31 | 1.33 | 22.8 | 5019 | target missed |
Note. : orbital semimajor axis; : orbital eccentricity; : orbital inclination; , : diameters of the primary and secondary components; : mutual separation distance; : visible spectral slope; : observation start time; : heliocentric distance at time of observation; : distance from JWST at time of observation; : phase angle of observation; : visible apparent magnitude at time of observation; : total exposure time across all four dithers. The orbital parameters and viewing geometry are computed by JPL Horizons. The visible spectral slope measurements are tabulated in fraser2021. The sizes and mutual separation of the binary components are taken from nesvorny2022.
@@ 2.1 Observation Design
The observations were carried out using the integral field unit (IFU) on the Near-Infrared Spectrograph (NIRSpec; jakobsen2022; boker2023). A set of four dithered exposures was collected for each target using the PRISM grating, which covers the 0.6–5.3 m wavelength range at an average spectral resolution of . To reduce detector read noise, all exposures were read out with the NRSIRS2RAPID method (moseley2010; rauscher2012). Due to the large uncertainties in the binary orbit solutions for our targets, no timing constraints were stipulated to schedule the observations to coincide with maximum component separation.
@@ 2.2 Target Selection
When selecting targets for observation as part of our Cycle 3 JWST program, we consulted the list of cold classicals with visible spectral slope measurements published in fraser2021. Restricting our search to confirmed binaries and adopting the spectral slope value of nm as the threshold dividing the less-red and red color subpopulations, we chose the three brightest blue binaries according to their minimum apparent magnitude during the Cycle 3 observing windows. The selected targets were 2001 XR254, 2003 HG57, and 2016 BP81, which have measured spectral slopes of , , and , respectively, in units of %/100 nm. Table 1 provides additional information about the targets, including component diameters, mutual separations, and the details of our observations.
@@ 2.3 Data Reduction
Data processing and spectral extraction were handled by the jwstspec pipeline (jwstspec)—a dedicated tool for customized end-to-end reduction of JWST spectroscopic datasets that was developed primarily for observations of small bodies in the solar system. This pipeline has been employed in numerous previous analyses of NIRSpec observations (e.g., emery2024; grundy2024; wong2024), and a full description of the overall pipeline workflow can be found in those publications. Beginning with the uncalibrated data, jwstspec passed through the first two stages of Version 1.17.1 of the official JWST pipeline (jwst), with necessary reference files drawn from context jwst_1322.pmap of the JWST Calibration Reference Data System. All pipeline parameters were left at their default values, except for the nsclean step in the second stage, which was manually activated to correct for residual read noise on the detector. The result of the pipeline processing was bias-corrected, flat-fielded, spatially rectified, wavelength-calibrated, and flux-calibrated IFU data cubes for each dithered exposure.
@@
The jwstspec pipeline offers several different spectral extraction techniques. For the results presented in this paper, we utilized the empirical PSF fitting method, which is optimized for faint point-source targets and yields substantial improvements in precision over standard circular aperture extraction (e.g., wong2024; pinillaalonso2025). In short, a moving median window of a specified width is applied along the wavelength axis to generate background-subtracted template PSFs, which are then normalized and fit to the central wavelength slices within the windows to obtain the best-fit multiplicative scaling factors for the templates. The fluxes are then measured from the scaled template PSFs within a user-defined circular aperture. See the referenced papers for full details on this method. For our observations, we set the background region as all pixels outside of a box centered on the measured centroid. The moving median window width was set to 51 (i.e., from to along the wavelength axis), though adjusting this value by up to did not substantively change the resultant spectra. Uncertainties on the fluxes were estimated by weighting the error arrays with the peak-normalized PSF template model and adding in quadrature the uncertainties of all pixels within the extraction aperture.
@@
The spectra extracted from the set of four dithered exposures were passed through a 21 pixel-wide moving median filter to remove outliers. The combined irradiance spectrum of each target was obtained by a simple mean of the four individual dithered spectra. Due to low instrument throughput and poor data quality at the extreme ends of the wavelength range, we trimmed the spectra to 0.7–5.1 m.
@@ 2.4 Accounting for Sensitivity Variations
We used NIRSpec observations of the G-type solar analog star SNAP-2 when deriving the reflectance spectra of our targets. By dividing the irradiance spectrum of an object by that of the star, processed in an analogous manner and extracted from an aperture with the same equivalent encircled flux, we self-consistently corrected for the wavelength-dependent flux losses outside of the extraction apertures while also removing any common-mode instrumental systematics that are present in the data. In the case of the blended binary 2001 XR254, our choice of the 4-pixel-radius aperture is roughly equivalent to two 3-pixel-radius apertures centered on the two components in terms of encircled flux, and we therefore used a 3-pixel-radius aperture when extracting the stellar spectrum. For 2016 BP81, the stellar spectrum was extracted with the same 2-pixel-radius aperture as the target spectrum.
The offset of 2016 BP81 relative to the nominal pointing region in the center of the FOV is almost , which is larger than the range subtended by the 4-point dither pattern. This raises the possibility of systematic sensitivity variations across the FOV that could bias the shape of the resultant reflectance spectrum when incorporating stellar observations in which the source was centered in a different region of the FOV. To investigate this issue, we considered two observations of SNAP-2, one that employed a 20-point dither pattern (Program #1128, Observation 2) and another that used the 4-point nod pattern (Program #4498, Observation 19).333For more on NIRSpec IFU dither and nod patterns, see https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-operations/nirspec-dithers-and-nods/nirspec-ifu-dither-and-nod-patterns. In the former case, the spread in pointing positions is similar to the 4-point dither pattern that was selected for our KBO targets, while in the latter, the internod separations are much larger, with the star placed at the vertices of a box centered on the middle of the FOV.
Figure 2 shows the irradiance spectra of SNAP-2 from the two observations, both extracted using a 3-pixel-radius aperture. Also plotted is the CALSPEC model spectrum of SNAP-2 (bohlin2014). The two NIRSpec spectra have largely equivalent shapes, except at wavelengths shorter than about 1.4 m: here, the spectrum measured from the 4-point nod observation drops off significantly, indicating a relative decrease in detector sensitivity toward the edges of the FOV. Meanwhile, the spectrum derived from the 20-point dither pattern is in good agreement with the CALSPEC model. It follows that the choice of one or the other SNAP-2 spectrum affects the continuum shape of the resultant reflectance spectrum, specifically the short-wavelength spectral slope.